OpenStudy (chaotic_butterflies):

Simplify (1 − sin x)(1 + sin x).

OpenStudy (misty1212):

HI!!

OpenStudy (chaotic_butterflies):

Hello

OpenStudy (misty1212):

multiply out you get \[1-\sin^2(x)\]

OpenStudy (misty1212):

then use a famous identity to rewrite it

OpenStudy (chaotic_butterflies):

Oh wow the first step was simple enough, but I'm not sure what you're talking about by a certain "famous identity"

OpenStudy (misty1212):

the mother of all trig identities is \[\cos^2(x)+\sin^2(x)=1\] so \[1-\sin^2(x)=\cos^2(x)\]

geerky42 (geerky42):

Aka Pythagorean Identity.

OpenStudy (chaotic_butterflies):

Okay, now I got it. Thank you very much! ^_^